OpenStudy (anonymous):

Show that 2^n+1 is 0(2^n)

OpenStudy (anonymous):

are you sure of what you just whrite

OpenStudy (anonymous):

That is the same question I asked the professor!

OpenStudy (anonymous):

Is that expression multiplied by zero ? 0(2^n)

geerky42 (geerky42):

No it's big O notation.

geerky42 (geerky42):

\(2^n+1 = O(2^n)\)

geerky42 (geerky42):

Right? @kattck

OpenStudy (anonymous):

That is what I was thinking at first but how it is written I have asked and it does represent the number 0

OpenStudy (anonymous):

Well, if it is, I haven't seen anything like this before. For me, that expression is equal to 0, there is no such 'n' for 2^2 + 1 to be zero.

OpenStudy (anonymous):

sorry, I was going to write 2^n+1

geerky42 (geerky42):

Well, it's not what big O means.

geerky42 (geerky42):

What class are you in? @kattck

OpenStudy (anonymous):

Yah this would be false statement correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

This is discrete Mathematic Comp Science

geerky42 (geerky42):

then I'm pretty sure it's big O.

OpenStudy (anonymous):

What subjects did your proffessor explained you so far?

geerky42 (geerky42):

Do you have note about it? What is this "0"?

geerky42 (geerky42):

What is definition of 0( f(x) )?

OpenStudy (anonymous):

I found a whole section on Big-"OH" notation!

OpenStudy (anonymous):

geerky42 You were right!

geerky42 (geerky42):

Yeah :) What is definition of big O, if you were given?

geerky42 (geerky42):

Then from here, we can figure out how to show etc

geerky42 (geerky42):

Or you can use this: https://www.cs.utexas.edu/users/tandy/big-oh.pdf

OpenStudy (anonymous):

ok so big OH is measuerment time of the execution of the algorithm

geerky42 (geerky42):

Well, yeah that's what big O is often used for.

geerky42 (geerky42):

In your question, by "show," you mean like prove?

OpenStudy (anonymous):

Yah this section is mostly about proving the algorithms.

geerky42 (geerky42):

OK, so from definition in link, \(2^n+1 = O(2^n)\) if there exist positive constant \(C_1\) and \(C_2\) such that \(2^n+1\le C_1~ 2^n\) for all \(n>C_2\) You get the idea, right?